Answer the following questions, which relate to the aggregate expenditures model: Instructions: Enter your answer as a whole number. a. Given the following: Ca = $130, Ig = $60, Xn = − $10, and G = $40, what is the economy’s equilibrium GDP? b. If real GDP in an economy is currently $250, will the economy’s real GDP rise, fall, or stay the same? c. Suppose that full-employment (and full-capacity) output in an economy is $250. If Ca = $180, Ig = $60, Xn = − $10, and G = $40, what will be the macroeconomic result?
Answer #1

(a) At equilibrium level , Y= Ca + Ig + G + Xn

Y = 130 + 60 + 40 + (-10)

Y = $220.
The economy's equilibrium GDP is $220.

(b) It is given that the real GDP of the economy is $250 and the current equilibrium level of GDP is $220. This means there must be unplanned inventory and the firms respond to it by building inventory and reducing output.

Thus the economy's real GDP will fall.

(c) AE = Ca + Ig + G + Xn

AE = 180 + 60 + 40 + (-10)

AE = $270.

It is given that the full employment output is $250 and the aggregate expenditure is $270. SInce, the AE is greater than the full employment level of output, there is an inflationary gap in the economy.